# Upper Approximation Bounds for Neural Oscillators

### v1

@@ Appendix C Explanation of the time dependence of $\Psi\left({\mathbf{x}},{\mathbf{0}},t\right)$
$\begin{array}[]{l}\left|{\mathbf{z}}_{2}\left(t,{\mathbf{z}}_{0}\right)-{\mathbf{z}}_{1}\left(t,{\mathbf{z}}_{0}\right)\right|\\ \;\;\;=\left|{\int_{0}^{t}{{{{\bf{z}}}_{2}^{\prime}}(\tau){\rm{d}}\tau}-\int_{0}^{t}{{{{\bf{z}}}_{1}^{\prime}}(\tau){\rm{d}}\tau}}\right|\\ \;\;\;=\left|{\int_{0}^{t}{{{\bf{z}}_{2,2}}(\tau){\rm{d}}\tau}-\int_{0}^{t}{{{\bf{z}}_{1,2}}(\tau){\rm{d}}\tau}}\right|+\left|{\int_{0}^{t}{{g_{1}}\left[{{{\bf{z}}_{1,1}}(\tau),{{\bf{z}}_{1,2}}(\tau),{\bf{u}}(\tau)}\right]}{\rm{d}}\tau-\int_{0}^{t}{{g_{2}}\left[{{{\bf{z}}_{2,1}}(\tau),{{\bf{z}}_{2,2}}(\tau),{\bf{u}}(\tau)}\right]{\rm{d}}\tau}}\right|\\ \;\;\;\leq\int_{0}^{t}{\left|{{{\bf{z}}_{2,2}}(\tau)-{{\bf{z}}_{1,2}}(\tau)}\right|{\rm{d}}\tau}+\int_{0}^{t}{\left|{{g_{1}}\left[{{{\bf{z}}_{1,1}}(\tau),{{\bf{z}}_{1,2}}(\tau),{\bf{u}}(\tau)}\right]-{g_{2}}\left[{{{\bf{z}}_{2,1}}(\tau),{{\bf{z}}_{2,2}}(\tau),{\bf{u}}(\tau)}\right]}\right|}{\rm{d}}\tau\\ \;\;\;\leq\int_{0}^{t}{\left|{{{\bf{z}}_{2,2}}(\tau)-{{\bf{z}}_{1,2}}(\tau)}\right|{\rm{d}}\tau}+\int_{0}^{t}{\left|{{g_{1}}\left[{{{\bf{z}}_{1,1}}(\tau),{{\bf{z}}_{1,2}}(\tau),{\bf{u}}(\tau)}\right]-{g_{1}}\left[{{{\bf{z}}_{2,1}}(\tau),{{\bf{z}}_{1,2}}(\tau),{\bf{u}}(\tau)}\right]}\right|}{\rm{d}}\tau\\ \;\;\;\;\;\;\;+\int_{0}^{t}{\left|{{g_{1}}\left[{{{\bf{z}}_{2,1}}(\tau),{{\bf{z}}_{1,2}}(\tau),{\bf{u}}(\tau)}\right]-{g_{1}}\left[{{{\bf{z}}_{2,1}}(\tau),{{\bf{z}}_{2,2}}(\tau),{\bf{u}}(\tau)}\right]}\right|}{\rm{d}}\tau\\ \;\;\;\;\;\;\;+\int_{0}^{t}{\left|{{g_{1}}\left[{{{\bf{z}}_{2,1}}(\tau),{{\bf{z}}_{2,2}}(\tau),{\bf{u}}(\tau)}\right]-{g_{2}}\left[{{{\bf{z}}_{2,1}}(\tau),{{\bf{z}}_{2,2}}(\tau),{\bf{u}}(\tau)}\right]}\right|}{\rm{d}}\tau.\end{array}$ (114)
@@
$\begin{array}[]{l}\left|{\mathbf{z}}_{2}\left(t,{\mathbf{z}}_{0}\right)-{\mathbf{z}}_{1}\left(t,{\mathbf{z}}_{0}\right)\right|\\ \;\;\;\leq\int_{0}^{t}{\left|{{{\bf{z}}_{2,2}}(\tau)-{{\bf{z}}_{1,2}}(\tau)}\right|{\rm{d}}\tau}+{L_{{g_{1}},1}}\int_{0}^{t}{\left|{{{\bf{z}}_{2,1}}(\tau)-{{\bf{z}}_{1,1}}(\tau)}\right|}{\rm{d}}\tau+{L_{{g_{1}},2}}\int_{0}^{t}{\left|{{{\bf{z}}_{1,2}}(\tau)-{{\bf{z}}_{2,2}}(\tau)}\right|}{\rm{d}}\tau\\ \;\;\;\;\;\;\;+\int_{0}^{t}{\left|{{g_{1}}\left[{{{\bf{z}}_{2,1}}(\tau),{{\bf{z}}_{2,2}}(\tau),{\bf{u}}(\tau)}\right]-{g_{2}}\left[{{{\bf{z}}_{2,1}}(\tau),{{\bf{z}}_{2,2}}(\tau),{\bf{u}}(\tau)}\right]}\right|}{\rm{d}}\tau\\ \;\;\;\leq{L_{{g_{1}}}}\int_{0}^{t}{\left|{{{\bf{z}}_{1}}(\tau)-{{\bf{z}}_{2}}(\tau)}\right|}{\rm{d}}\tau+\mathop{\max}\limits_{\tau\in[0,T]}\left|{{g_{1}}\left[{{{\bf{z}}_{2,1}}(\tau),{{\bf{z}}_{2,2}}(\tau),{\bf{u}}(\tau)}\right]-{g_{2}}\left[{{{\bf{z}}_{2,1}}(\tau),{{\bf{z}}_{2,2}}(\tau),{\bf{u}}(\tau)}\right]}\right|T,\\ \;\;\;\mathop{\leq}\limits^{(a)}{e^{T{L_{{g_{1}}}}}}T\mathop{\max}\limits_{\tau\in[0,T]}\left|{{g_{1}}\left[{{{\bf{z}}_{2,1}}(\tau),{{\bf{z}}_{2,2}}(\tau),{\bf{u}}(\tau)}\right]-{g_{2}}\left[{{{\bf{z}}_{2,1}}(\tau),{{\bf{z}}_{2,2}}(\tau),{\bf{u}}(\tau)}\right]}\right|,\end{array}$ (115)
@@ Appendix E Approach for Training the Neural Oscillator
where $\mathbf{z}(0)=\mathbf{0}$, $\mathbf{y}(0)={\it{\Pi}}\left[\mathbf{z}_{1}(0),\mathbf{u}(0),t\right]$, $\Delta{t_{i}}=t_{i+1}-t_{i}$, $0=t_{0}<t_{1}<t_{2}<\dots<t_{I}=T$. Given available output function samples $\hat{\mathbf{y}}_{l}(t_{i})=\left[\hat{y}_{l,1}(t_{i}),\hat{y}_{l,2}(t_{i}),\dots,\hat{y}_{l,q}\right]^{\top}$ caused by the input samples $\mathbf{u}_{l}(t_{i})$, $l=1,2,\dots,L$, the $r$-th power loss function $\ell_{r}$ is used to learn the trainable parameters of the MLPs ${\it{\Gamma}}(\cdot)$ and ${\it{\Pi}}(\cdot)$ in Eq. (116),
$\ell_{r}=\frac{1}{LIq}{\sum\limits_{l=1}^{L}{\sum\limits_{i=0}^{I}\sum\limits_{j=1}^{q}{{\left|{y_{l,j}}(t_{i})-\hat{y}_{l,j}(t_{i})\right|}^{r}}}},$ (118)

### v2

@@ Appendix C Explanation of the time dependence of $\Psi\left({\mathbf{x}},{\mathbf{0}},t\right)$
$\left|{\mathbf{z}}_{2}\left(t,{\mathbf{z}}_{0}\right)-{\mathbf{z}}_{1}\left(t,{\mathbf{z}}_{0}\right)\right|$ (118)
$\;\;\;=\left|{\int_{0}^{t}{{{{\bf{z}}}_{2}^{\prime}}(\tau){\rm{d}}\tau}-\int_{0}^{t}{{{{\bf{z}}}_{1}^{\prime}}(\tau){\rm{d}}\tau}}\right|$
$\;\;\;=\left|{\int_{0}^{t}{{{\bf{z}}_{2,2}}(\tau){\rm{d}}\tau}-\int_{0}^{t}{{{\bf{z}}_{1,2}}(\tau){\rm{d}}\tau}}\right|+\left|{\int_{0}^{t}{{g_{1}}\left[{{{\bf{z}}_{1,1}}(\tau),{{\bf{z}}_{1,2}}(\tau),{\bf{u}}(\tau)}\right]}{\rm{d}}\tau-\int_{0}^{t}{{g_{2}}\left[{{{\bf{z}}_{2,1}}(\tau),{{\bf{z}}_{2,2}}(\tau),{\bf{u}}(\tau)}\right]{\rm{d}}\tau}}\right|$
$\;\;\;\leq\int_{0}^{t}{\left|{{{\bf{z}}_{2,2}}(\tau)-{{\bf{z}}_{1,2}}(\tau)}\right|{\rm{d}}\tau}+\int_{0}^{t}{\left|{{g_{1}}\left[{{{\bf{z}}_{1,1}}(\tau),{{\bf{z}}_{1,2}}(\tau),{\bf{u}}(\tau)}\right]-{g_{2}}\left[{{{\bf{z}}_{2,1}}(\tau),{{\bf{z}}_{2,2}}(\tau),{\bf{u}}(\tau)}\right]}\right|}{\rm{d}}\tau$
$\;\;\;\leq\int_{0}^{t}{\left|{{{\bf{z}}_{2,2}}(\tau)-{{\bf{z}}_{1,2}}(\tau)}\right|{\rm{d}}\tau}+\int_{0}^{t}{\left|{{g_{1}}\left[{{{\bf{z}}_{1,1}}(\tau),{{\bf{z}}_{1,2}}(\tau),{\bf{u}}(\tau)}\right]-{g_{1}}\left[{{{\bf{z}}_{2,1}}(\tau),{{\bf{z}}_{1,2}}(\tau),{\bf{u}}(\tau)}\right]}\right|}{\rm{d}}\tau$
$\;\;\;\;\;\;\;+\int_{0}^{t}{\left|{{g_{1}}\left[{{{\bf{z}}_{2,1}}(\tau),{{\bf{z}}_{1,2}}(\tau),{\bf{u}}(\tau)}\right]-{g_{1}}\left[{{{\bf{z}}_{2,1}}(\tau),{{\bf{z}}_{2,2}}(\tau),{\bf{u}}(\tau)}\right]}\right|}{\rm{d}}\tau$
$\;\;\;\;\;\;\;+\int_{0}^{t}{\left|{{g_{1}}\left[{{{\bf{z}}_{2,1}}(\tau),{{\bf{z}}_{2,2}}(\tau),{\bf{u}}(\tau)}\right]-{g_{2}}\left[{{{\bf{z}}_{2,1}}(\tau),{{\bf{z}}_{2,2}}(\tau),{\bf{u}}(\tau)}\right]}\right|}{\rm{d}}\tau.$
@@
$\left|{\mathbf{z}}_{2}\left(t,{\mathbf{z}}_{0}\right)-{\mathbf{z}}_{1}\left(t,{\mathbf{z}}_{0}\right)\right|$ (119)
$\;\;\;\leq\int_{0}^{t}{\left|{{{\bf{z}}_{2,2}}(\tau)-{{\bf{z}}_{1,2}}(\tau)}\right|{\rm{d}}\tau}+{L_{{g_{1}},1}}\int_{0}^{t}{\left|{{{\bf{z}}_{2,1}}(\tau)-{{\bf{z}}_{1,1}}(\tau)}\right|}{\rm{d}}\tau+{L_{{g_{1}},2}}\int_{0}^{t}{\left|{{{\bf{z}}_{1,2}}(\tau)-{{\bf{z}}_{2,2}}(\tau)}\right|}{\rm{d}}\tau$
$\;\;\;\;\;\;\;+\int_{0}^{t}{\left|{{g_{1}}\left[{{{\bf{z}}_{2,1}}(\tau),{{\bf{z}}_{2,2}}(\tau),{\bf{u}}(\tau)}\right]-{g_{2}}\left[{{{\bf{z}}_{2,1}}(\tau),{{\bf{z}}_{2,2}}(\tau),{\bf{u}}(\tau)}\right]}\right|}{\rm{d}}\tau$
$\;\;\;\leq{L_{{g_{1}}}}\int_{0}^{t}{\left|{{{\bf{z}}_{1}}(\tau)-{{\bf{z}}_{2}}(\tau)}\right|}{\rm{d}}\tau+\mathop{\max}\limits_{\tau\in[0,T]}\left|{{g_{1}}\left[{{{\bf{z}}_{2,1}}(\tau),{{\bf{z}}_{2,2}}(\tau),{\bf{u}}(\tau)}\right]-{g_{2}}\left[{{{\bf{z}}_{2,1}}(\tau),{{\bf{z}}_{2,2}}(\tau),{\bf{u}}(\tau)}\right]}\right|T,$
$\;\;\;\mathop{\leq}\limits^{(a)}{e^{T{L_{{g_{1}}}}}}T\mathop{\max}\limits_{\tau\in[0,T]}\left|{{g_{1}}\left[{{{\bf{z}}_{2,1}}(\tau),{{\bf{z}}_{2,2}}(\tau),{\bf{u}}(\tau)}\right]-{g_{2}}\left[{{{\bf{z}}_{2,1}}(\tau),{{\bf{z}}_{2,2}}(\tau),{\bf{u}}(\tau)}\right]}\right|,$
@@ Appendix E Approach for Training the Neural Oscillator
where $\mathbf{z}(0)=\mathbf{0}$, $\mathbf{y}(0)={\it{\Pi}}\left[\mathbf{z}_{1}(0),\mathbf{u}(0),t\right]$, $\Delta{t_{i}}=t_{i+1}-t_{i}$, $0=t_{0}<t_{1}<t_{2}<\dots<t_{I}=T$. Given available output function samples $\hat{\mathbf{y}}_{l}(t_{i})=\left[\hat{y}_{l,1}(t_{i}),\hat{y}_{l,2}(t_{i}),\dots,\hat{y}_{l,q}(t_{i})\right]^{\top}$ caused by the input samples $\mathbf{u}_{l}(t_{i})$, $l=1,2,\dots,L$, the $r^{\text{th}}$ power loss function $\ell_{r}$ is used to learn the trainable parameters of the MLPs ${\it{\Gamma}}(\cdot)$ and ${\it{\Pi}}(\cdot)$ in Eq. (120),
$\ell_{r}=\frac{1}{LIq}{\sum\limits_{l=1}^{L}{\sum\limits_{i=0}^{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}I-1}\sum\limits_{j=1}^{q}{{\left|{y_{l,j}}(t_{i})-\hat{y}_{l,j}(t_{i})\right|}^{r}}}},$ (122)